the mean of 6 numbers is 24.If one number is excluded, the mean of remaining numbers becomes 22.find the excluded number.
Answer by Princess F(1)

x/6 =24
x = 24X6= 144
x =144/6= 24
Mean =24
Remaining numbers=5x22=110
Average number=144-110=34
6 numbers could be 22+22+22+22+22+34÷6 =24
If one number is excluded
22+22+22+22+22÷5 = 110/5 =22
Excluded number= 34
Ans= 34